OpenStudy (anonymous):

find the critical points of f on the domain or given interval \[(e ^{x}+e ^{-x})/2\]

OpenStudy (zehanz):

You'll have to look for the derivative, and solve: ((e^x+e^-x)/2)' = 0

OpenStudy (anonymous):

the derivative is the exact same thing as the original in this case

OpenStudy (anonymous):

but I'm not sure how they got 0 as the answer

OpenStudy (zehanz):

It's not exactly the same: (e^x)' = e^x, but (e^-x)' = -e^-x

OpenStudy (zehanz):

So, (the factor two is not important) you end up with e^x - e^-x=0 Or: e^x=e^-x, which has only 0 as a solution.

OpenStudy (anonymous):

oh wow stupid mistake, my fault. sorry about that lol

OpenStudy (zehanz):

No problem, happens to all of us from time to time...

OpenStudy (anonymous):

A critical point is where the graph of the function turns. This is where the first derivative is zero. So for, \[y =\frac{ ( e^x +e ^{-x}) }{ 2 }\] So, \[\frac{ d }{ dx } (\frac{ ( e^x +e ^{-x}) }{ 2 }) = \frac{ 1 }{ 2 } (e^x - e^{-x} )\] When does this equal zero? Well, it equals zero when, \[\frac{ 1 }{ 2 } (e^{x} - e^{-x}) = 0 \] \[e^{x} - e^{-x}=0\] \[e^{x}=e^{-x}\] Taking the natural logarithm on both sides, we have \[x = -x\] And so, \[x=0\]